OpenStudy (anonymous):

help? 1/2< or equal to x+1/3 < 3/4

OpenStudy (shadowlegendx.):

\[\frac{ 1 }{ 2 } \le x+ \frac{ 1 }{ 3 } < \frac{ 3 }{ 4 }\]

OpenStudy (shadowlegendx.):

You need to find the LCD, Least Common Denominator to solve this

OpenStudy (shadowlegendx.):

Once you do, convert the fractions and solve